Foster's Repair ShopFoster's Repair Shop

has a monthly target operating income of

$50,000.

Variable expenses are

75​%

of​ sales, and monthly fixed expenses are

$11,000.

Read the requirements

1.

Compute the monthly margin of safety in dollars if the shop achieves its income goal.

2.

Express

FosterFoster​'s

margin of safety as a percentage of target sales.

3.

What is

FosterFoster​'s

operating leverage factor at the target level of operating​ income?

4.

Assume that the company reaches its target. By what percentage will the​ company's operating income fall if sales volume declines by

10​%?
Answer #1

1. Contribution margin ratio = 100%-75% = 25%
Sales to achieve target profit = (Fixed Cost + Target Income) / Contribution margin ratio
= ($11000+50000)/25% = $244000

Break Even sales = Fixed Cost / Contribution margin ratio
= $11000 / 25% = $44000

Margin of Safety = Sales - Breakeven Sales
= $244000 - 44000 = $200000

2. Margin of Safety (%) = Margin of Safety / Target Sales x 100
= $200000 / 244000 x 100 = 81.97%

3. Operating Leverage = Contribution Margin / Net Income
= ($244000x25%) / $50000 = 1.22

4. Operating income will fall by 1.22 x 10% i.e. 12.20%